Hello, my name is kalea I need help whether a given ordered pair is a solution of an equation in two variables. The homework problem I need help with is:
Y=4x - 1
There are 3 solutions to these it can either be
(2,5) (0,-1) (-1,6) (-2, -9) (1,3)
I tried doing This
Y= 4 x - 1 (2, 5)
5+4(2)-1
and I got
3=1 which did not work out and i used the same strategies for the following and it was not working out.

Answer by jim_thompson5910(35256)

Let's try the point (2,5)
x = 2, y = 5

y = 4x - 1
y = 4(2) - 1 ... replace x with 2
5 = 4(2) - 1 ... replace y with 5
5 = 8 - 1
5 = 7 ... FALSE

the equation is false when (x,y) = (2,5). So (2,5) is NOT a solution to the given equation.

-------------------------------------------------
Let's try the point (0,-1)
x = 0, y = -1

y = 4x - 1
y = 4(0) - 1 ... replace x with 0
-1 = 4(0) - 1 ... replace y with -1
-1 = 0 - 1
-1 = -1 ... TRUE

the equation is true when (x,y) = (0,-1). So (0,-1) is definitely a solution to the given equation.

-------------------------------------------------
Let's try the point (-1,6)
x = -1, y = 6

y = 4x - 1
y = 4(-1) - 1 ... replace x with -1
6 = 4(-1) - 1 ... replace y with 6
6 = -4 - 1
6 = -5 ... FALSE

the equation is false when (x,y) = (-1,6). So (-1,6) is NOT a solution to the given equation.

-------------------------------------------------
Let's try the point (-2,-9)
x = -2, y = -9

y = 4x - 1
y = 4(-2) - 1 ... replace x with -2
-9 = 4(-2) - 1 ... replace y with -9
-9 = -8 - 1
-9 = -9 ... TRUE

the equation is true when (x,y) = (-2,-9). So (-2,-9) is definitely a solution to the given equation.

-------------------------------------------------
Let's try the point (1,3)
x = 1, y = 3

y = 4x - 1
y = 4(1) - 1 ... replace x with 1
3 = 4(1) - 1 ... replace y with 3
3 = 4 - 1
3 = 3 ... TRUE

the equation is true when (x,y) = (1,3). So (1,3) is definitely a solution to the given equation.

-------------------------------------------------
The three solutions are:
(0,-1)
(-2,-9)
(1,3)